OpenStudy (anonymous):

Ted is not particularly creative. He uses the pickup line "If I could arrange the alphabet, I'd put U and I together." The random variable x is the number of girls Ted approaches before encountering one who reacts positively. Determine whether the the table describes a probability distribution. If it does, find its mean and standard deviation. x P(x) 1 0.001 2 0.024 3 0.131 4 0.236 5 0.335

OpenStudy (anonymous):

if the numbers add up to one, the answer is yes

OpenStudy (anonymous):

if \[\sum_{i=1}^{n} P(x_{i})=1\] then it is a probability distribution

OpenStudy (anonymous):

in your case n=5

OpenStudy (anonymous):

Look at the probabilities, firstly all them should add upto 1

OpenStudy (anonymous):

clearly n=5 as you have 5 outcomes, \[\sum_{i=1}^{5}P(x_{i})=P(x_{1})+P(x_{2})+P(x_{3})+P(x_{4})+P(x_{5})\]\[\sum_{i=1}^{5}P(x_{i})=0.001+0.024+0.131+0.236+0.335\]

OpenStudy (anonymous):

@slomomo check if that sum is equal to 1

OpenStudy (anonymous):

The sum does not equal 1, it equals 0.727, so does this mean it is not a probability distribution?

OpenStudy (anonymous):

@slomomo correct! It's not a probability distribution

OpenStudy (anonymous):

So would that just be the answer or is it another kind of distribution? Because in that question, which I did not include, it said "Determine whether the table describes a probability solution." So as I stated I am taking a pre-test and we have to show our work so I guess I would say that this is not a probability distribution table and type in the set of numbers to show they do not equal 1. Would that be correct?

OpenStudy (anonymous):

I meant distribution, not solution. Sorry!

OpenStudy (anonymous):

@Nishant_Garg I posted above that in that pre-test question, it said "Determine whether the table describes a probability distribution" so since the answer is no, I was just wondering is it considered any other probability table or should I just select the NO box and then we have a box that pops up to show our work so I guess I will just show that those do not add up to 1 but 0.727. Does that seem correct or would there be a formula for this?

OpenStudy (anonymous):

Since it's not a probability distribution, you don't need to find its means and standard deviation, and the question ends here